# Continuous-variable quantum key distribution system: A review and perspective

### v2

@@ ABSTRACT
Quantum key distribution provides secure keys with information-theoretic security ensured by the principle of quantum mechanics. The continuous-variable version of quantum key distribution using coherent states offers the advantages of its compatibility with telecom industry, e.g., using commercial laser and homodyne detector, is now going through a booming period. In this review article, we describe the principle of continuous-variable quantum key distribution system, focus on protocols based on coherent states, whose systems are gradually moving from proof-of-principle lab demonstrations to in-field implementations and technological prototypes. We start by reviewing the theoretical protocols and the current security status of these protocols. Then, we discuss the system structure, the key module, and the mainstream system implementations. The advanced progress for future applications are discussed, including the digital techniques, system on chip and point-to-multipoint system. Finally, we discuss the practical security of the system and conclude with promising perspectives in this research field.
@@ I INTRODUCTION
Since 1984, quantum key distribution (QKD) Bennett and Brassard (1984) has ushered in an era of secure communications using quantum methods by providing information-theoretic secure key distribution Ekert (1991); Gisin et al. (2002); Scarani et al. (2009); Pirandola et al. (2020); Xu et al. (2020); Cao et al. (2022); Portmann and Renner (2022). The combination of this method with one-time-pad encryption provides the ultimate protection for the transmission of confidential messages. In general, for simplified implementations, QKD protocols are formulated in a prepare-and-measure fashion, where classical information is encoded in non-orthogonal quantum states: these are randomly prepared by Alice (the sender) and then transmitted to Bob (the receiver) through an insecure quantum channel. At the output of the channel, the states are measured by Bob to retrieve the encoded classical information. The quantum no-cloning theorem dictates that an unknown quantum state cannot be reliably cloned Wootters, Wootters, and Zurek (1982), ensuring long-term security based on physical principles Diamanti et al. (2016) against unlimited computational power.
@@
So far, various QKD protocols with discrete variables have been proposed to support the long-distance and practical-secure system implementations Bennett and Brassard (1984); Ekert (1991); Lütkenhaus (2000); Inoue, Waks, and Yamamoto (2002); Hwang (2003); Scarani et al. (2004); Barrett, Hardy, and Kent (2005); Stucki et al. (2005); Wang (2005); Lo, Ma, and Chen (2005); Acín et al. (2007); Inamori, Lütkenhaus, and Mayers (2007); Pironio et al. (2009); Masanes, Pironio, and Acín (2011); Braunstein and Pirandola (2012); Lo, Curty, and Qi (2012); Sasaki, Yamamoto, and Koashi (2014); Lucamarini et al. (2018); Arnon-Friedman et al. (2018); Vazirani and Vidick (2019), including the decoy states experiments Zhao et al. (2006); Peng et al. (2007); Rosenberg et al. (2007); Schmitt-Manderbach et al. (2007); Yuan, Sharpe, and Shields (2007); Yin et al. (2008); Wang et al. (2008); Dixon et al. (2008); Rosenberg et al. (2009); Yuan et al. (2009); Liu et al. (2010); Chen et al. (2010); Wang et al. (2013a); Boaron et al. (2018), the measurement-device-independent (MDI) experiments Rubenok et al. (2013); Liu et al. (2013); Da Silva et al. (2013); Tang et al. (2014a, b, c); Wang et al. (2015a); Valivarthi et al. (2015); Yin et al. (2016); Tang et al. (2016a, b); Comandar et al. (2016); Kaneda et al. (2017); Wang et al. (2017a); Valivarthi et al. (2017); Liu et al. (2018a); Wei et al. (2020), the twin-field experiments Minder et al. (2019); Wang et al. (2019a); Liu et al. (2019); Zhong et al. (2019); Fang et al. (2020); Chen et al. (2020); Pittaluga et al. (2020); Chen et al. (2021a); Wang et al. (2022a); Liu et al. (2023), the system on chip Ma et al. (2016a); Sibson et al. (2017a, b); Bunandar et al. (2018); Ding et al. (2017); Paraïso et al. (2019); Wei et al. (2020), and so on Lee et al. (2014); Guan et al. (2015); Takesue et al. (2015); Wang et al. (2015b); Zhong et al. (2015); Korzh et al. (2015); Mirhosseini et al. (2015); Sit et al. (2017); Li et al. (2016a); Yuan et al. (2016); Islam et al. (2017). Specifically, the twin-field QKD with a 3-station scheme has significantly promoted the development of the long-distance QKD, where the total distance can break the PLOB bound Pirandola et al. (2017). These achievements have resulted in the long-haul point-to-point connection up to 1000 km Liu et al. (2023), high-speed metropolitan system Tanaka et al. (2012); Lucamarini et al. (2013); Fröhlich et al. (2017); Yuan et al. (2018); Li et al. (2023a), and field deployed QKD network Elliott et al. (2005); Peev et al. (2009); Chen et al. (2009); Sasaki et al. (2011); Chen et al. (2021b), even with satellite-to-ground links Yin et al. (2017a); Liao et al. (2017); Ren et al. (2017); Yin et al. (2017b); Liao et al. (2018a); Chen et al. (2021c). Furthermore, it is worth noting a distinct category of protocols where information is encoded on the quadrature of light that is continuous-variable. The use of such continuous-variable quantum information carriers, instead of qubits, constitutes a potent and alternative approach for QKD Ralph (1999); Cerf, Lévy, and Assche (2001); Grosshans and Grangier (2002); Weedbrook et al. (2004); Braunstein and van Loock (2005); Wang et al. (2007); Weedbrook et al. (2012); Lam and Ralph (2013); Diamanti and Leverrier (2015) and more broadly, for quantum information processing Braunstein and van Loock (2005); Cerf, Leuchs, and Polzik (2007); Furusawa and Takei (2007); Andersen, Leuchs, and Silberhorn (2010); Weedbrook et al. (2012); Van Meter (2014); Adesso, Ragy, and Lee (2014); Andersen et al. (2015); Kurizki et al. (2015); Serafini (2017).
@@
Continuous-variable QKD (CV-QKD) using coherent states Grosshans and Grangier (2002); Weedbrook et al. (2004) is now currently experiencing a booming period due to its compatibility with telecom industry, e.g., using commercial continuous-wave laser and coherent receiver. This potential has led to significant advancements in CV-QKD, including protocol design, security analysis, and system implementation (See Fig. 1). The security of CV-QKD protocol using Gaussian-modulated coherent states was initially proved under asymptotic conditions García-Patrón and Cerf (2006); Navascués, Grosshans, and Acín (2006), and later extended to the finite-size regime with universal composability against collective attack Leverrier (2015), and general attacks by exploiting Gaussian de Finetti theorem Leverrier (2017). In addition to continuous modulation, the discrete modulation of coherent states has also been well investigated Li, Zhang, and Guo (2018); Ghorai et al. (2019); Lin, Upadhyaya, and Lütkenhaus (2019); Denys, Brown, and Leverrier (2021); Matsuura et al. (2021). Along with the improving security proof, the implementation of CV-QKD system has progressed from the initial proof-of-principle demonstration to the second stage with swift advancements in high performance and system robustness. Currently, it has entered the third stage where a modern architecture is evolving with the benefit of being fully compatible with coherent communication.
@@
Figure 1: Selected key developments in CV-QKD. On both sides of the arrow are theoretical research progress (left) and experimental research progress (right). Here, the theroetical research progress includes the protocol proposal, security proof and optimization of protocol steps. The experimental progress includes the development of in-line LO and local LO systems.
@@
Figure 2: The overview of the CV-QKD system, including the typical systems, key techniques and the appliactions. Here we mainly show the long-distance, chip-based and high-speed system, as well as the key techniques including the reconciliation, chip-based devices and digital techniques.
@@
During the initial phases of the CV-QKD system, the primary challenge was to overcome the 3 dB limit, which was solved by the reverse reconciliation Grosshans et al. (2003a). Subsequently, the CV-QKD system progressed towards allowing long-distance transmission to facilitate two-user interconnection across a wide range without a trusted relay. Developments of the reconciliation resulted in enhanced error correction capability even with an extremely low signal-to-noise ratio (SNR) Leverrier et al. (2008), which played a significant role in the long-distance system covering a distance from 25 km Lodewyck et al. (2007) up to 80 km Jouguet et al. (2013). At this stage, both the quantum signal and local oscillator (LO) are generated by the same laser of transmitter and co-propagated in the quantum channel, known as the in-line LO system, which contributes to the suppression of phase noise when the quantum and LO interfere for coherent detection. However, a significant challenge towards achieving long-distance and stable transmission is to reduce crosstalk between the strong LO and the weak quantum signal. The most effective current approach is using pulsed signals with high extinction ratio, then combining polarization multiplexing and time-division multiplexing Jouguet et al. (2012a, 2013); Huang et al. (2016a, b); Zhang et al. (2019a, b, 2020a). This methodology has resulted in the longest lab experiment over 202 km Zhang et al. (2020a), the longest field test of 50 km Zhang et al. (2019a), the long-term test of a 3-node CV-QKD network in Qingdao, China Zhang et al. (2020b), and the first chip-based system Zhang et al. (2019b).
@@
In 2015, an alternative scheme was proposed, which relaxed the requirement of the extremely high isolation by generating LO inside the receiver and using a pilot assisted phase recovery to suppress the phase noise introduced by the different laser source Qi et al. (2015); Soh et al. (2015). As the pilot signal has significantly lower power compared to the LO, the high-extinction pulse generation for time division multiplexing is no longer required. Instead, frequency-division multiplexing is widely used, which simplifies the signal generation, and sometimes can be combined with polarization multiplexing for better isolation. For higher secret key rate and better phase recovery, the repetition rate of the system is gradually enhanced, where digital techniques in classical optical communication systems, such as pulse shaping and matched filter, are introduced to overcome the limited bandwidth of devices. Further, more and more digital algorithms for a CV-QKD system are developed, including the de-multiplexing, impairment compensation, synchronization etc., which then drive the innovation in system architecture. To date, the local LO scheme contributes to the high-speed system with the repetition rate of 5 GBaud, resulting in 190 Mbps secret key rate at 5 km Wang et al. (2022b), as well as a flexible network deployment, which has been demonstrated by the software-defined CV-QKD network in Madrid, Spain Aguado et al. (2019).
@@
In all this panorama, the present review aims at providing an overview of the most important results and the most recent advances in the field of CV-QKD system. After a brief introduction of the general notions, we review the main CV-QKD protocols and security analysis in Sec. II. The system structure and key modules are reviewed in Sec. III, and the typical currently-achievable implementations are detailed in Sec. IV, including the in-line LO systems, local LO systems, systems co-existing with classical networks, and so on. We then discuss the advanced progress for future applications, such as digital continuous-variable system and point-to-multipoint network, in Sec. V. Finally, we will discuss the practical security of the system in Sec. VI and conclude with promising perspectives in this research field in Sec. VII.
@@ II.1 Basic notions of continuous-variable systems
A continuous-variable system is a typical infinite dimensional quantum system. Suppose that a CV system consists of a sequence of $n$ modes in the Hilbert space $\mathcal{H}=\otimes_{i=1}^{n}\mathcal{H}_{i}$, there are $n$ pairs of annihilation and creation operators $\left\{\hat{a}_{i},\hat{a}^{{\dagger}}_{i}\right\}$ with i=1,2,,ni=1,2,...,n, which satisfies
@@ II.2 A historical outline of CV-QKD protocols and the current security status
The first CV-QKD protocol was proposed in 1999, using squeezed states to achieve secret key distribution Ralph (1999). However, due to the challenges in preparing squeezed states, a protocol of using coherent states and homodyne detection to distribute secret key was proposed in 2002 Grosshans and Grangier (2002), namely the GG02 protocol. Because the coherent states can be easily generated by a laser, this protocol has received an increasing attention in recent years. Subsequently the protocol based on Gaussian modulated coherent states got further developments. The no-switching protocol was reported in 2004 Weedbrook et al. (2004), in which heterodyne detection instead of homodyne detection was used. In 2009, heterodyne detection was also utilized in the Gaussian modulated squeezed-state protocol and an improvement of performance was found García-Patrón and Cerf (2009).
Although Gaussian modulated CV-QKD protocols have undergone extensive study and development, there remain technical challenges to implementing ideal Gaussian modulation in experiments. Therefore, discrete modulated CV-QKD protocol Leverrier and Grangier (2009) was proposed. The initial discrete modulated CV protocol is the four-state modulation protocol, where a modulation method similar to QPSK in classical communications was used. Soon afterwards, A. Leverrier Leverrier and Grangier (2011) highlighted that CV protocols can be used for secret key distribution with multi-dimensional discrete modulation. In addition to discrete modulation, other modulation methods are also considered, including unidimensional modulation Usenko and Grosshans (2015). It simplifies the modulation process at the Alice side can be compared with the GG02 protocol when the excess noise is small. Most of the theoretical analysis of the CV-QKD protocols are based on fiber channels, and recently, the study is extended to the free space scenario for considerations of satellite quantum communications. Pirandola (2021c)
Generally, the classification of standard one-way CV-QKD protocols, in which the quantum state passes through a single channel, can follow the type of the used quantum states (coherent or squeezed), the methods of modulation (Gaussian modulation, unidimensional modulation or discrete modulation) or the type of measurement (homodyne or heterodyne). In addition to the one-way protocols, various other protocols correspond to different application scenarios are proposed, such as two-way protocols, source-device-independent protocols, measurement-device-independent protocols, and so forth Goncharov et al. (2022).

### v3

# Continuous-variable quantum key distribution system: past, present, and future

@@ ABSTRACT
Quantum key distribution provides secure keys with information-theoretic security ensured by the principle of quantum mechanics. The continuous-variable version of quantum key distribution using coherent states offers the advantages of its compatibility with telecom industry, e.g., using commercial laser and homodyne detector, is now going through a booming period. In this review article, we describe the principle of continuous-variable quantum key distribution system, focus on protocols based on coherent states, whose systems are gradually moving from proof-of-principle lab demonstrations to in-field implementations and technological prototypes. We start by reviewing the theoretical protocols and the current security status of these protocols. Then, we discuss the system structure, the key module, and the mainstream system implementations. The advanced progresses for future applications are discussed, including the digital techniques, system on chip and point-to-multipoint system. Finally, we discuss the practical security of the system and conclude with promising perspectives in this research field.
@@ I INTRODUCTION
Since 1984, quantum key distribution (QKD) Bennett and Brassard (1984) has ushered in an era of secure communications using quantum methods by providing information-theoretic secure key distribution Ekert (1991); Gisin et al. (2002); Scarani et al. (2009); Pirandola et al. (2020); Xu et al. (2020); Cao et al. (2022); Portmann and Renner (2022). The combination of this method with one-time-pad encryption provides the ultimate protection for the transmission of confidential messages. In general, for simplified implementations, QKD protocols are formulated in a prepare-and-measure fashion, where the classical information is encoded on non-orthogonal quantum states: they are randomly prepared by Alice (the sender) and then transmitted to Bob (the receiver) through an insecure quantum channel. At the output of the channel, the states are measured by Bob to retrieve the encoded classical information. The quantum no-cloning theorem dictates that an unknown quantum state cannot be reliably cloned Wootters, Wootters, and Zurek (1982), ensuring long-term security based on physical principles Diamanti et al. (2016) against unlimited computational power.
@@
So far, various QKD protocols with discrete variables have been proposed to support the long-distance and practical-secure system implementations Bennett and Brassard (1984); Ekert (1991); Lütkenhaus (2000); Inoue, Waks, and Yamamoto (2002); Hwang (2003); Scarani et al. (2004); Barrett, Hardy, and Kent (2005); Stucki et al. (2005); Wang (2005); Lo, Ma, and Chen (2005); Acín et al. (2007); Inamori, Lütkenhaus, and Mayers (2007); Pironio et al. (2009); Masanes, Pironio, and Acín (2011); Braunstein and Pirandola (2012); Lo, Curty, and Qi (2012); Sasaki, Yamamoto, and Koashi (2014); Lucamarini et al. (2018); Arnon-Friedman et al. (2018); Vazirani and Vidick (2019), including the decoy states experiments Zhao et al. (2006); Peng et al. (2007); Rosenberg et al. (2007); Schmitt-Manderbach et al. (2007); Yuan, Sharpe, and Shields (2007); Yin et al. (2008); Wang et al. (2008); Dixon et al. (2008); Rosenberg et al. (2009); Yuan et al. (2009); Liu et al. (2010); Chen et al. (2010); Wang et al. (2013a); Boaron et al. (2018), the measurement-device-independent (MDI) experiments Rubenok et al. (2013); Liu et al. (2013); Da Silva et al. (2013); Tang et al. (2014a, b, c); Wang et al. (2015a); Valivarthi et al. (2015); Yin et al. (2016); Tang et al. (2016a, b); Comandar et al. (2016); Kaneda et al. (2017); Wang et al. (2017a); Valivarthi et al. (2017); Liu et al. (2018a); Wei et al. (2020), the twin-field experiments Minder et al. (2019); Wang et al. (2019a); Liu et al. (2019); Zhong et al. (2019); Fang et al. (2020); Chen et al. (2020); Pittaluga et al. (2020); Chen et al. (2021a); Wang et al. (2022a); Liu et al. (2023), the system on chip Ma et al. (2016a); Sibson et al. (2017a, b); Bunandar et al. (2018); Ding et al. (2017); Paraïso et al. (2019); Wei et al. (2020), and so on Lee et al. (2014); Guan et al. (2015); Takesue et al. (2015); Wang et al. (2015b); Zhong et al. (2015); Korzh et al. (2015); Mirhosseini et al. (2015); Sit et al. (2017); Li et al. (2016a); Yuan et al. (2016); Islam et al. (2017). Specifically, the twin-field QKD with a 3-station scheme has significantly promoted the development of the long-distance QKD, where the total distance can break the PLOB bound Pirandola et al. (2017). These achievements have resulted in the long-haul point-to-point connection up to 1000 km Liu et al. (2023), high-speed metropolitan system Tanaka et al. (2012); Lucamarini et al. (2013); Fröhlich et al. (2017); Yuan et al. (2018); Li et al. (2023a), and field deployed QKD network Elliott et al. (2005); Peev et al. (2009); Chen et al. (2009); Sasaki et al. (2011); Chen et al. (2021b), even with satellite-to-ground links Yin et al. (2017a); Liao et al. (2017); Ren et al. (2017); Yin et al. (2017b); Liao et al. (2018a); Chen et al. (2021c). Furthermore, it is worth noting a distinct category of protocols where the information is encoded on the quadrature of light that is continuous-variable. The use of such continuous-variable quantum information carriers, instead of qubits, constitutes a potent and alternative approach for QKD Ralph (1999); Cerf, Lévy, and Assche (2001); Grosshans and Grangier (2002); Weedbrook et al. (2004); Braunstein and van Loock (2005); Wang et al. (2007); Weedbrook et al. (2012); Lam and Ralph (2013); Diamanti and Leverrier (2015) and more broadly, for quantum information processing Braunstein and van Loock (2005); Cerf, Leuchs, and Polzik (2007); Furusawa and Takei (2007); Andersen, Leuchs, and Silberhorn (2010); Weedbrook et al. (2012); Van Meter (2014); Adesso, Ragy, and Lee (2014); Andersen et al. (2015); Kurizki et al. (2015); Serafini (2017).
@@
Continuous-variable QKD (CV-QKD) using coherent states Grosshans and Grangier (2002); Weedbrook et al. (2004) is now currently experiencing a booming period due to its compatibility with telecom industry, e.g., using commercial continuous-wave laser and coherent receiver. This potential has led to significant advancements in CV-QKD, including the protocol design, security analysis, and system implementation (See Fig. 1). The security of CV-QKD protocol using Gaussian-modulated coherent states was initially proved under asymptotic conditions García-Patrón and Cerf (2006); Navascués, Grosshans, and Acín (2006), and later extended to the finite-size regime with universal composability against collective attack Leverrier (2015), and general attacks by exploiting Gaussian de Finetti theorem Leverrier (2017). In addition to continuous modulation, the discrete modulation of coherent states has also been well investigated Li, Zhang, and Guo (2018); Ghorai et al. (2019); Lin, Upadhyaya, and Lütkenhaus (2019); Denys, Brown, and Leverrier (2021); Matsuura et al. (2021). Along with the improving security proof, the implementation of CV-QKD system has progressed from the initial proof-of-principle demonstration to the second stage with swift advancements in high performance and system robustness. Currently, it has entered the third stage where a modern architecture is evolving with the benefit of being fully compatible with classical optical coherent communication.
@@
Figure 1: Selected key developments in CV-QKD. On both sides of the arrow are theoretical research progresses (left) and experimental research progresses (right). Here, the theoretical research progresses includes the protocol proposal, security proof and optimization of protocol steps. The experimental progresses includes the development of in-line LO and local LO systems.
@@
Figure 2: The overview of the CV-QKD system, including the typical systems, key techniques and the applications. Here we mainly show the long-distance, chip-based and high-speed system, as well as the key techniques including the reconciliation, chip-based devices and digital techniques.
@@
During the initial phase of the CV-QKD system, the primary challenge was to overcome the 3 dB limit, which was solved by the reverse reconciliation Grosshans et al. (2003a). Subsequently, the CV-QKD system progressed towards allowing long-distance transmission to facilitate two-user interconnection across a wide range without a trusted relay. Developments of the reconciliation resulted in enhanced error correction capability even with an extremely low signal-to-noise ratio (SNR) Leverrier et al. (2008), which played a significant role in the long-distance system covering a distance from 25 km Lodewyck et al. (2007) up to 80 km Jouguet et al. (2013). At this stage, both the quantum signal and local oscillator (LO) are generated by the same laser of transmitter and co-propagated in the quantum channel, known as the in-line LO system, which contributes to the suppression of phase noise when the quantum signal and LO interfere for coherent detection. However, a significant challenge towards achieving long-distance and stable transmission is to reduce crosstalk between the strong LO and the weak quantum signal. The most effective current approach is using pulsed signals with high extinction ratio, then combining polarization multiplexing and time-division multiplexing Jouguet et al. (2012a, 2013); Huang et al. (2016a, b); Zhang et al. (2019a, b, 2020a). This methodology has resulted in the longest lab experiment over 202 km Zhang et al. (2020a), the longest field test of 50 km Zhang et al. (2019a), the long-term test of a 3-node CV-QKD network in Qingdao, China Zhang et al. (2020b), and the first chip-based system Zhang et al. (2019b).
@@
In 2015, an alternative scheme was proposed, which relaxed the requirement of the extremely high isolation by generating LO inside the receiver and using a pilot assisted phase recovery to suppress the phase noise introduced by the different laser source Qi et al. (2015); Soh et al. (2015). As the pilot signal has significantly lower power compared to the LO, the high-extinction pulse generation for time-division multiplexing is no longer required. Instead, frequency-division multiplexing is widely used, which simplifies the signal generation, and sometimes can be combined with polarization multiplexing for better isolation. For higher secret key rate and better phase recovery, the repetition rate of the system is gradually enhanced, where digital techniques in classical optical communication systems, such as the pulse shaping and matched filter, are introduced to overcome the limited bandwidth of devices. Further, more and more digital algorithms for a CV-QKD system are developed, including the de-multiplexing, impairment compensation, synchronization etc., which then drive the innovation in system architecture. To date, the local LO scheme contributes to the high-speed system with the repetition rate of 5 GBaud, resulting in 190 Mbps secret key rate at 5 km Wang et al. (2022b), as well as a flexible network deployment, which has been demonstrated by the software-defined CV-QKD network in Madrid, Spain Aguado et al. (2019).
@@
In all this panorama, the present review aims at providing an overview of the most important results and the most recent advances in the field of CV-QKD system. After a brief introduction of the general notions, we review the main CV-QKD protocols and security analysis in Sec. II. The system structure and key modules are reviewed in Sec. III, and the typical currently-achievable implementations are detailed in Sec. IV, including the in-line LO systems, local LO systems, systems co-existing with classical networks, and so on. We then discuss the advanced progress for future applications, such as the digital continuous-variable system and point-to-multipoint network, in Sec. V. Finally, we will discuss the practical security of the system in Sec. VI and conclude with promising perspectives in this research field in Sec. VII.
@@ II.1 Basic notions of continuous-variable systems
The continuous-variable system is a typical infinite dimensional quantum system. Suppose that a CV system consists of a sequence of $n$ modes in the Hilbert space $\mathcal{H}=\otimes_{i=1}^{n}\mathcal{H}_{i}$, there are $n$ pairs of annihilation and creation operators $\left\{\hat{a}_{i},\hat{a}^{{\dagger}}_{i}\right\}$ with i=1,2,,ni=1,2,...,n, which satisfies
@@ II.2 A historical outline of CV-QKD protocols and the current security status
The first CV-QKD protocol was proposed in 1999, using squeezed states to achieve secret key distribution Ralph (1999). However, due to the challenges in preparing squeezed states, a protocol of using coherent states and homodyne detection to distribute secret key was proposed in 2002 Grosshans and Grangier (2002), namely the GG02 protocol. Because the coherent states can be easily generated by a laser, this protocol has received an increasing attention in recent years. Subsequently the protocol based on Gaussian modulated coherent states got further developments. The no-switching protocol was reported in 2004 Weedbrook et al. (2004), in which the heterodyne detection instead of homodyne detection was used. In 2009, heterodyne detection was also utilized in the Gaussian modulated squeezed-state protocol and an improvement of performance was found García-Patrón and Cerf (2009).
Although Gaussian modulated CV-QKD protocols have undergone extensive studies and developments, there remain technical challenges to implementing ideal Gaussian modulation in experiments. Therefore, discrete modulated CV-QKD protocol Leverrier and Grangier (2009) was proposed. The initial discrete modulated CV protocol is the four-state modulation protocol, where a modulation method similar to QPSK in classical communications was used. Soon afterwards, A. Leverrier Leverrier and Grangier (2011) highlighted that CV protocols can be used for secret key distribution with multi-dimensional discrete modulation. In addition to discrete modulation, other modulation methods are also considered, including unidimensional modulation Usenko and Grosshans (2015). It simplifies the modulation process at the Alice side can be compared with the GG02 protocol when the excess noise is small. Recently, a phase-sensitive multimode protocol is proposed, which achieves higher secret key rate and better excess noise tolerance Su et al. (2023). Most of the theoretical analysis of the CV-QKD protocols are based on fiber channels, and recently, the study is extended to the free space scenario for considerations of satellite quantum communications. Pirandola (2021c)
Generally, the classification of standard one-way CV-QKD protocols, in which the quantum state passes through a single channel, can follow the type of the used quantum states (coherent or squeezed), the methods of modulation (Gaussian modulation, unidimensional modulation or discrete modulation) or the type of measurement (homodyne or heterodyne). In addition to the one-way protocols, various other protocols correspond to different application scenarios are proposed, such as two-way protocols, source-device-independent protocols, MDI protocols, and so forth Goncharov et al. (2022).